OpenStudy (anonymous):

Which literal equation is equivalent to 1.4 = kx A. k= x divided by 1.4 B. x= 1.4 x K C. k= 1.4 divided by x D. x= k divided by 1.4

OpenStudy (jonnyvonny):

c

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your Welcome

OpenStudy (anonymous):

can you help with this one too. Solve for r. A. r = Ka/p B. r = Kap C. r = Kp/a D. r = a/Kp

OpenStudy (anonymous):

I think its c for the top one :D